Solve -4(5-2n)=8(-6-5n)
Answer by rmromero(383)


Solve -4(5-2n)=8(-6-5n) Use distributive property
-4(5) -4(-2n) = 8(-6) +8(-5n)
-20 + 8n = -48 - 40n Simplify.subtract 8n both sides.
Add 48 both sides.
48n = -28 Divide both sides by 48
28
n = - ___ Reduce to lowest term
48

7
n = - ____
12
Checking:
-4(5-2n) = 8(-6-5n)
-4(5-2(-7/12))=8(-6-5(-7/12))
-4(5+7/6) = 8(-6+35/12)
-20 + 28/6 = -48 + 70/3
-120 - 28 -288 + 140
___________ = __________
6 6
-148 -148
_______ = _________ ---->> true
6 6

I hope this helps
Visit my wesbite http://onlinemathtutorial.tripod.com
Sign my guestbook. thanks